The rectangular plate shown above has a width of 5cm and length of 4cm at temperature at 15 Celsius. The plate is made of metal with the thermal expansion coefficient a=7.0*10^-6 1/Celsius. What is the final area of the plate when the temperature increases from T=15 Celsius to T=72 Celsius

state your answe in cm2 to the nearest 0.0001 of the cm2


1
Expert's answer
2021-09-30T10:20:22-0400

The arear of the hot plate is


"A_\text{hot}=a(1+\alpha\Delta T)·b(1+\alpha\Delta T)=ab(1+\alpha \Delta T)^2=\\=20.0160\text{ cm}^2."